# Inspection-Guided Randomization: A Flexible and Transparent Restricted Randomization Framework for Better Experimental Design

### v1

@@ Appendix C Vignette 2 Data-generating Process
$(\sigma_{individual,j},\sigma_{school,j})=(0.1\cdot sd(\mu_{\cdot j}),0.2\cdot sd(\mu_{\cdot j}))$; $\gamma\in\{0.25,0.50,0.75\}$; ($\theta^{w},\theta^{b1},\theta^{b2})=(0.2,0.1,0.01)$; $\boldsymbol{\beta}=\begin{pmatrix}1&1\end{pmatrix}^{T}$; $q=0.25$; and $\tau=\delta=0.3\cdot SD(\mathbf{X}\boldsymbol{\beta})$. For IGR, we enumerate the pool of candidate allocations using a complete cluster randomization design, then apply a thresholding decision rule. In the main text, we show results for an enumerated pool of size $M=100,000$ and an accepted set of size $m=500$. Results for $M\in\{5000,1000,25000,50000,1000000,200000\}$ and $m\in\{100,500,1000\}$ are in Section D.
Appendix D Additional Results from Vignette 2
Figure D.1: [Vignette 2, sparsity $\gamma=0.5$, $N=4000$] Visualizations for the Evaluation & Adaptation step when $f^{(\texttt{I+M})}$ is used as the fitness function. (top row) $m^{(\texttt{InvMinEuclideanDist})}$ is only takes on a handful of distinct values, so discriminatory power is driven by $m^{(\texttt{MaxMahalanobis})}$. (middle row) While $m^{(\texttt{InMinEuclideanDist})}$ is coarser than $m^{(\texttt{FracCtrlExposed})}$, it still helps rule out the worst allocations. (bottom row) Pairwise assignment correlation under IGR and IGRg is comparable to that of CR.
Bias %$\%\text{Bias}_{CR}$ Var %$\%\text{Var}_{CR}$ Rej. Rate Design Fitness CR -0.12$\pm$0.05 Ref 0.48$\pm$0.07 Ref 0.24$\pm$0.03 IGR 0.25*MaxMahalanobis + 0.75*FracExpo 85.53$\pm$1.85 82.33$\pm$23.82 0.32$\pm$0.15 0.50*MaxMahalanobis + 0.50*FracExpo 90.62$\pm$2.42 17.01$\pm$2.85 0.89$\pm$0.06 0.75*MaxMahalanobis + 0.25*FracExpo 96.00$\pm$2.06 4.90$\pm$0.25 1.00$\pm$0.00 0.25*MaxMahalanobis + 0.75*InvMinEuclidDist 95.68$\pm$1.24 32.96$\pm$19.87 0.66$\pm$0.30 0.50*MaxMahalanobis + 0.50*InvMinEuclidDist 97.45$\pm$1.43 14.96$\pm$7.74 0.90$\pm$0.12 0.75*MaxMahalanobis + 0.25*InvMinEuclidDist 98.68$\pm$1.73 7.83$\pm$2.32 1.00$\pm$0.01 IGRg 0.25*MaxMahalanobis + 0.75*FracExpo 84.63$\pm$2.70 62.68$\pm$10.89 0.35$\pm$0.07 0.50*MaxMahalanobis + 0.50*FracExpo 90.46$\pm$3.16 9.70$\pm$1.17 0.98$\pm$0.01 0.75*MaxMahalanobis + 0.25*FracExpo 96.41$\pm$2.25 2.09$\pm$0.27 1.00$\pm$0.00 0.25*MaxMahalanobis + 0.75*InvMinEuclidDist 94.62$\pm$0.92 25.24$\pm$17.51 0.74$\pm$0.24 0.50*MaxMahalanobis + 0.50*InvMinEuclidDist 97.40$\pm$1.65 6.84$\pm$4.40 0.98$\pm$0.03 0.75*MaxMahalanobis + 0.25*InvMinEuclidDist 98.87$\pm$1.17 2.75$\pm$0.73 1.00$\pm$0.00
Table D.1: [Vignette 2, sparsity $\gamma=0.75$, $N=4000$] Bias, sample variance, and rejection rate of effect estimates under IGR compared to complete cluster randomization (CR). Applying greater weight to the interference-related inspection metric is more effective for bias reduction. Applying greater weight to the balance-related inspection metric is more effective for reducing variance. The coarser interference metric $m^{(\texttt{InvMinEuclideanDist})}$ helps to decrease interference bias, albeit not as effectively $m^{(\texttt{FracExpo})}$.
Figure D.2: [Vignette 2, sparsity $\gamma=0.5$, $N=4000$, $w_{balance}=0.5$, $w_{interference}=0.5$] Bias, variance, and rejection rate versus number of enumerated candidate allocations $M$ and the number of accepted allocations $m$. Thin lines represent results from a single data sample. Thick lines represent means taken across five data samples. When $m^{(\texttt{FracExpo})}$ is used as the interference metric, increasing $M$ results in greater bias and variance reductions relative to complete randomization (CR). When $m^{(\texttt{InvMinEuclidDist})}$ is used as the interference metric, increasing $M$ results in a slightly greater bias reduction and a more noticeable gain in variance reduction. Rejection rate increases as $M$ increases for both interference metrics.
Figure D.3: [Vignette 2, sparsity $\gamma=0.5$, $N=4000$] Bias, variance, and rejection rate versus magnitude of the weight applied to the balance metric, $m^{(\texttt{MaxMahalanobis})}$. Across plotted points, the weight applied to the interference metric is equal to 1 - $w_{\text{MaxMahalanobis}}$. As $w_{\text{MaxMahalanobis}}$ increases, IGR yields less bias reduction compared to complete randomization (CR). IGR can increase bias compared to CR when $w_{\text{MaxMahalanobis}}$ is large ($\gtrapprox 0.8$ for $m^{(\texttt{FracExpo})}$ and $\gtrapprox 0.4$ for $m^{(\texttt{InvMinEuclidDist})}$). As $w_{\text{MaxMahalanobis}}$ increases, IGR yields greater variance reduction compared to CR. IGR can increase variance compared to CR when $w_{\text{MaxMahalanobis}}$ is small ($\lessapprox 3$ for $m^{(\texttt{FracExpo})}$). Rejection rate increases as $w_{\text{MaxMahalanobis}}$ increases.
Appendix E Vignette 0: Multi-arm Experiment
In this additional vignette, we illustrate how to use IGR in the design of a multi-arm experiment, in a manner similar to Ciolino et al., (2019), Watson et al., (2021), and Zhou et al., (2011). In a multi-arm experiment, we simultaneously evaluate multiple interventions under a single trial protocol. To estimate treatment effects, we make pairwise comparisons between arms, often between each intervention arm and a single control arm. The design desideratum is to allocate treatments such that covariates are balanced across arms. We first discuss the choices we make in the Specification and Enumeration steps of IGR, then present results from a simulated experiment of a multi-arm classroom-based intervention.
E.1 Specification and Enumeration for designing a multi-arm experiment
Specification
We assume there is one control arm, and the remaining arms are treated with different interventions. We define the target effect estimands to be the average treatment effects comparing each of the treatments to the control, $\tau_{h}=\frac{1}{N}\sum_{i=1}^{N}Y_{i}(h)-Y_{i}(0)$ for all $h=1,\dots k-1$, where we assume arm 00 is the control arm and arms $11,\dots k-$ 1 are treated.
We assess covariate balance using the metrics $m^{(\texttt{MaxMahalanobis})}$ and $m^{(\texttt{SumMaxAbsSMD})}$, which are defined similarly to $m^{(\texttt{MaxMahalanobis-G})}$ and $m^{(\texttt{SumMaxAbsSMD-G})}$ (4.1.2) except there is no salient attribute that is excluded. Since there is only a single desideratum in this vignette, the aggregator function $\Theta$ is simply the identity function, and the fitness function is the inspection metric itself. Lastly, we use the thresholding restriction rule as defined in Vignette 1 (4.1).
Enumeration
To enumerate candidate allocations, we sample from a complete randomization assignment mechanism that assigns an equal number of individuals to each arm.
E.2 Simulated multi-arm experiment of educational feedback types
Data Generation
We use a similar data-generating process as the one detailed in Section B to generate baseline covariates, where all covariates are drawn in the same manner except for gender. In this vignette, we sample gender as $male_{i}\sim\text{Bern}(0.7)$ instead of setting a fixed gender composition in the study population. We again designate exam grade the outcome variable and define an outcome-generation process for the observed outcome $Y_{i}^{obs}$ for individual $i$ under treatment $z_{i}$ as
$Y_{i}^{obs}=\text{exam}_{i}+\sum_{j=1}^{k}\tau_{j}\mathbbm{1}\{z_{i}=j\}$ (E.1)
where $\{\tau_{j}\}_{j=1}^{k}\in\mathbb{R}^{k}$ are the treatment effects. The different arms correspond to different types of feedback on class assignments.
[…]

### v2

@@ Appendix C Data Generation for Experiment With Interference
$(\sigma_{individual,j},\sigma_{school,j})=(0.1\cdot sd(\mu_{\cdot j}),0.2\cdot sd(\mu_{\cdot j}))$; $\gamma=0.50$; ($\theta^{w},\theta^{b1},\theta^{b2})=(0.2,0.1,0.01)$; $\boldsymbol{\beta}=\begin{pmatrix}1&1\end{pmatrix}^{T}$; $q=0.25$; and $\tau=\delta=0.3\cdot SD(\mathbf{X}\boldsymbol{\beta})$.
Appendix D Toy Example Visualizations for the Evaluation & Adaptation Step of IGR
When applying the Evaluation & Adaptation step of IGR to our simulated example (see Section 4.2), we used visualizations to help check for discriminatory power, desiderata trade-offs, and over-restriction. In this section, we provide a more detailed description and rationale for each type of visualization used. We use synthetic toy examples to assist in our explanations.
D.1 Discriminatory Power
We highlighted discriminatory power as one check to make during the Evaluation & Adaptation step of IGR. To perform this check, we can plot a histogram of the fitness scores across all candidate allocations. In Figure D.1, we show two example distributions of fitness scores that one might encounter. A distribution with poor or low discriminatory power is one where many of the allocations have the same or similar score, which is easily identifiable in a histogram. A distribution with high discriminatory power is one where the scores are more uniformly distributed, such that the better-scoring allocations are clearly distinguished from the worse-scoring allocations.
Figure D.1: Histograms of fitness function scores across candidate allocations for a fitness function with low discriminatory power (orange) and high discriminatory power (blue). Note that smaller values of $f(\mathbf{z})$ are considered more desirable.
D.2 Desiderata Trade-offs
We also highlighted desiderata trade-offs as an important check to make in the Evaluation & Adaptation step. When there are two desiderata – and two corresponding inspection metrics for measuring them – we can use two-dimensional histograms and scatterplots to help with performing this check.
In the two-dimensional histogram, we plot the distribution of pairs of inspection metric values across all candidate allocations. This helps to reveal the extent to which desiderata trade-offs exist in the candidate space. In Figures D.2 - D.4, we show three example histograms that suggest small, medium, and large trade-offs, respectively. When the trade-off between the metrics is non-existent or small, then smaller values of one metric also tend to correspond to smaller values of the other metric. On the other hand, for medium or large trade-offs, smaller values of one metric tend to correlate with larger values of the other metric.
In the scatterplot, we plot pairs of inspection metric values across the accepted allocations. This helps to reveal the extent to which the chosen fitness function and restriction rule “navigate” the trade-offs in the candidate allocation space. In Figures D.2 - D.4, we show accepted allocations under a fitness function that linearly combines the two inspection metrics. When there are medium to large trade-offs between desiderata, then the set of accepted allocations is more sensitive to the weights that are used in the linear combination.
Figure D.2: (left) Histogram of inspection metric values across candidate allocations. Notice that the two metrics are positively correlated, meaning smaller and better values of one metric correspond to smaller and better values of the other. (right) Scatterplots of inspection metric values across accepted allocations under fitness functions that apply different weights to the metrics. When the trade-off between the metrics is small, the choice of weighting scheme does not have a large effect on the set of accepted allocations.
Figure D.3: (left) Histogram of inspection metric values across candidate allocations. Notice that the two metrics are somewhat negatively correlated, meaning smaller and better values of one metric tend to correspond to larger and worse values of the other. (right) Scatterplots of inspection metric values across accepted allocations under fitness functions that apply different weights to the metrics. By applying a greater weight to Metric 1 ($w_{M1}=0.75,w_{M2}=0.25$), we prioritize accepting allocations with a smaller value of Metric 1, at the cost of permitting some allocations with a larger value of Metric 2. An analogous trade-off occurs when we apply greater weight to Metric 2 ($w_{M1}=0.25,w_{M2}=0.75$).
Figure D.4: (left) Histogram of inspection metric values across candidate allocations. Notice that there is a strong negative correlation between the two metrics, meaning the best values of one metric tend to correspond to the worst values of the other. (right) Scatterplots of inspection metric values across accepted allocations under fitness functions that apply different weights to the metrics. When the trade-off is large, we may not be able to avoid accepting some allocations with a small value of one metric and a large value of the other, no matter what kind of weighting scheme we use.
D.3 Over-restriction
Lastly, we highlight checking for over-restriction. In particular, we examine the extent to which there is pairwise assignment correlation across the accepted allocations. As mentioned in the Evaluation & Adaptation step in Section 4.2, pairwise assignment correlation can reduce power. As an estimate of pairwise correlation, for each pair of units $i,j$, we calculate the fraction of accepted allocations where $i,j$ are assigned to the same arm, $\hat{P}(z_{i}=z_{j})=\sum_{\mathbf{z}\in\mathcal{Z}_{pool},p_{\mathbf{z}}>0}\mathbf{1}\{z_{i}=z_{j}\}$. We then plot a histogram of these fractions. When there is high assignment correlation, then we expect the histogram to have mass at values closer to 0 and 1, which correspond to a pair $i,j$ never being assigned to the same arm and always being assigned to the same arm, respectively. When there is little to no assignment correlation, the histogram should be largely concentrated around 0.5. In Figure D.5, we show an example of low and high pairwise assignment correlation.
Figure D.5: Histograms of approximate pairwise assignment correlation to check for over-restriction. For high pairwise assignment correlation (blue), the histogram has mass closer to 0 and to 1, indicating that there are many pairs of units $i,j$ that are nearly always assigned to different or to the same arm. For low pairwise assignment correlation (orange), pairs of units are assigned to the same arm around half the time.
Appendix E IGR for a Simulated Group Formation Experiment
In this section, we present another example use case for IGR where prior restricted randomization strategies have not been applied.
E.1 Simulated Group Formation Experiment
Group Formation Experiments for Composition-Dependent Effects
In a group formation experiment, the challenge is to minimize differences in most baseline covariates, while creating large contrasts in a salient covariate of interest. Individuals are randomized into different groups, typically to construct groups of certain compositions that are believed to have different responses to the intervention (Basse et al.,, 2024). For example, a group may be the set of college freshmen assigned to the same dorm room, and composition may be the number of individuals in a room who were “high academic achievers” in high school (Sacerdote,, 2001). We consider group formation trials where composition is defined in terms of a categorical salient attribute (e.g., high or low prior academic achievement) and is count-based (e.g., number of individuals with high academic achievement).
We illustrate how to use IGR in the design of a group formation experiment where we are interested in evaluating how group composition influences the effect of a group-level intervention. Various kinds of inter-group comparisons may be of interest, including between treated and control groups with the same compositions (to target the treatment effect under a fixed composition) and between treated and treated groups with different compositions (to target the effect of the composition under a fixed treatment). In addition, we may also want to make difference-in-differences comparisons between the treatment effect under one composition and the treatment effect under a different composition.
The design goals in this setting are two-fold. First, we would like to sort individuals into groups such that certain pre-specified group compositions are realized. We focus in particular on the case where composition is characterized by the proportion of the group that possesses a particular salient attribute. Second, we would like to maintain balance on all covariates other than those used to define composition. A key point of this vignette is to highlight that in the Enumeration step of IGR, we can use different methods to sample allocations. Here, we sample from a blocking-based group formation assignment mechanism.
Simulation Set-up
We simulate the study data set to encode theory-based covariance between student characteristics (age, gender, and college major), latent traits (ability and confidence), and course performance (homework score and final exam grade). In particular, course performance is assumed to vary with students’ latent ability and confidence, which in turn depend on age, gender, and major. Ability is modeled to increase with age and is higher for some majors than other majors. In the context of STEM majors where the majority of students are men, we model confidence to be higher among men than women. We also model higher confidence for certain STEM-related majors.
Let $\mathbf{x}_{i}=\begin{pmatrix}age_{i}&gender_{i}&major_{i}&hw_{i}&exam_{i}\end{pmatrix}^{T}$ denote observed baseline covariates for individual $i$ and $\mathbf{u}_{i}=\begin{pmatrix}ability_{i}&confidence_{i}\end{pmatrix}^{T}$ denote latent, unobserved covariates. To generate $gender_{i}$, we randomly set half of the study sample to be men. Although having a fixed gender composition in the study population is not strictly necessary, it makes it easier to implement a group formation strategy that assembles groups with certain gender compositions. In a real-world setting, we could construct a study population with a fixed gender ratio by using an enrollment approach that intentionally targets this ratio. We use the following data-generating process for the remaining covariates:
$age_{i}$ $\sim\text{Unif}(19,25)$
$major_{i}$ $\sim\{1,2,3\}\text{ with probability }\{0.5,0.3,0.2\}$
$ability_{i}$ $=(age_{i}-\bar{age})/\sigma_{age}-0.5\cdot\mathbf{1}\{gender_{i}=1\}+\mathbf{1}\{major_{i}=1\}+\varepsilon$
$confidence_{i}$ $=\mathbf{1}\{gender_{i}=1\}+\mathbf{1}\{major_{i}=2\}+\varepsilon$
$hw_{i}$ $=ability_{i}+\epsilon$
$exam_{i}$ $=ability_{i}+confidence_{i}+\varepsilon$
$\varepsilon$ $\sim N(0,1)$
$Y_{i}^{obs}=\text{exam}_{i}+\mathbf{e}_{c(i)}^{T}\boldsymbol{\tau}z_{i}$ (E.1)
where $\mathbf{e}_{c(i)}\in\mathbb{R}^{l}$ is a vector that has a 1 in the $c(i)^{th}$ entry and a 0 everywhere else, and $\boldsymbol{\tau}\in\mathbb{R}^{l}$ is a vector of treatment effects. Entry $j$ in $\boldsymbol{\tau}$ can be interpreted as the effect of the intervention for the $j^{th}$ composition type. We run simulations with groups of composition types $\{c_{a}^{1},c_{a}^{2},c_{a}^{3}\}=\{0.5,0.3,0.7\}$ (that is, 50% men, 30% men, and 70% men), with 20 individuals per group.
E.2 Step-by-Step Application of IGR
E.2.1 Specification
We specify the target effect estimands as the average treatment effects under particular group compositions of interest. We specify one inspection metric that inspects allocations for balance and one for whether desired group compositions are realized. For balance, we extend the commonly used Mahalanobis distance metric to the group formation experiment setting. Let $\mathbf{z}\in\{0,1\}^{N_{G}}$ denote the vector of treatment assignments across groups. Since there are both random treatment assignments $\mathbf{z}$ and random group assignments $\mathbf{g}$, we use fitness functions of the form $f(\mathbf{z},\mathbf{g},\mathbf{X})$. We define $m^{(\texttt{Mahalanobis-G})}$ as
$m^{(\texttt{Mahalanobis-G})}(\mathbf{z},\mathbf{g},\mathbf{X})$ $=\max_{(h_{1},h_{2})\in\mathcal{G}\times\mathcal{G}}\left(\bar{\mathbf{X}}_{\setminus a}^{(h_{1})}-\bar{\mathbf{X}}_{\setminus a}^{(h_{2})}\right)S_{\setminus a}^{-1}\left(\bar{\mathbf{X}}_{\setminus a}^{(h_{1})}-\bar{\mathbf{X}}_{\setminus a}^{(h_{2})}\right)^{T}$ (E.2)
where $\mathcal{G}=\{0,1,2,\cdots,N_{G}-1\}$ is the set of group indices, $\bar{X}_{\cdot j}^{(g)}$ and $s_{\cdot j}^{(h)}$ are the mean and standard deviation of covariate $j$ in group $h$, $\mathbf{\bar{X}}_{\setminus a}^{(h)}$ is the vector of covariate means in group $h$ excluding the salient attribute $a$, and $S_{\setminus a}$ is the sample covariance matrix excluding $a$. To ensure allocations have the desired group compositions, we define the metric $m^{(\texttt{DesiredComps})}$, which takes on value 1 if all desired compositions are realized in at least two groups and is 0 otherwise. Combining the balance and composition realization metrics, we define the fitness function
$\begin{split}f^{(\texttt{Mahalanobis-GD})}(\mathbf{z},\mathbf{g},\mathbf{X})&=\begin{cases}m^{(\texttt{Mahalanobis-G})}(\mathbf{z},\mathbf{g},\mathbf{X})&m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X})=1\\ \infty&m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X})=0\end{cases}\end{split}$ (E.3)
E.2.2 Enumeration
For simplicity, we assume that we want to form groups of equal sizes such that there are exactly two groups with each composition, one of which is assigned treatment and one assigned control. To enumerate a candidate allocation, we first block on the salient attribute, then randomly sample subsets of individuals to place into groups such that the compositions of interest are achieved. For each pair of groups with the same composition, we randomly assign one to treatment and one to control. Under this group formation randomization assignment mechanism, $m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X})$ is guaranteed to equal 1. We refer to this assignment mechanism as group formation randomization (GFR). Note that, if we were to instead sample candidate allocations from an assignment mechanism that is agnostic to group composition, it would require enumerating an extremely large number of allocations to obtain any allocations that have the desired group compositions. Using the GFR assignment mechanism, we enumerate $M=100,000$ candidate allocations.
E.2.3 Restriction
We use the restriction rule $r(s)=0\text{ if }\ s\geq s^{*},r(s)=1\text{ if }s<s^{*}$, where $s^{*}$ is the $0.5^{\text{th}}$-percentile of scores for the enumerated candidate allocations. This restriction rule filters out all but the top $0.5^{\text{th}}$-percentile of the $100,000$ enumerated allocations, resulting in $m=500$ accepted allocations.
E.2.4 Evaluation & Adaptation
As with the other simulated experiment, we perform visual checks to evaluate the fitness function and restriction rule. As before, we use a histogram of scores to evaluate the discriminatory power of the fitness function and a histogram of pairwise assignment correlation to evaluate over-restriction. Because we have a single continuous inspection metric, we do not need to evaluate desiderata trade-offs. Instead, we want to ensure that we achieve good balance on each individual covariate, which is not necessarily guaranteed since we used a balance metric that averages across covariates. To check for per-covariate balance, we construct a boxplot of the standardized mean difference between treatment arms for each covariate.
Figure E.1: Visual checks for discriminatory power, over-restriction (via approximate pairwise assignment correlation), and balance under fitness function $f^{(\texttt{Mahalanobis-GD})}$. Note that abilityability and confidenceconfidence are latent attributes and were not included in the balance metric.
E.2.5 Pre-registration
We pre-register the set of 500 accepted allocations, each with equal probability $\frac{1}{500}$ of being drawn. Additionally, we pre-register the fitness function $f^{(\texttt{Mahalanobis-GD})}$ and the threshold restriction rule $r$.
E.2.6 Randomization
We sample one of the 500 accepted allocations uniformly at random.
E.3 How IGR Improves Effect Estimates
In Table E.1, we show that, as expected, IGR with both a balance and composition metric reduces the root mean-squared error (RMSE) of treatment effect estimates compared to a benchmark group formation randomization that uses a composition metric alone.
[…]
Table E.1: Root mean-squared error (RMSE) in treatment effect estimates and rejection rate of sharp null for IGR in a group formation experiment. IGR with both a balance and a composition realization metric reduces RMSE across group comparisons compared to a benchmark group formation randomization (GFR; equivalently, IGR with a composition metric alone). Means and standard deviations are computed over 5 data samples.
Appendix F IGR Bias Properties
In this section, we prove bias properties in the difference-in-means estimator applied to an IGR-designed experiment.
F.1 Unbiasedness Through the Mirror Property
Assumption 1 (Equal-sized Arms).
We assume that the design assigns the same number of individuals to the treated and control arms, so $\sum_{i=1}^{n}Z_{i}=\sum_{i=1}^{n}1-Z_{i}=\frac{n}{2}$.
Proof of Proposition 1.
The proof is identical to that of Morgan and Rubin, (2012) for unbiasedness in re-randomization. Because $\tilde{P}_{IGR}$ satisfies the mirror property, by definition, $\tilde{P}_{IGR}(\mathbf{Z}=\mathbf{z})=\tilde{P}_{IGR}(\mathbf{Z}=\mathbf{1}-\mathbf{z})$. By implication, $\tilde{P}_{IGR}(Z_{i}=0)=\tilde{P}_{IGR}(Z_{i}=1)=\frac{1}{2}$. The difference-in-means estimator is $\hat{\tau}:=\frac{\sum_{i=1}^{n}Y_{i}Z_{i}}{\sum_{i=1}^{n}Z_{i}}-\frac{\sum_{i=1}^{n}Y_{i}(1-Z_{i})}{\sum_{i=1}^{n}1-Z_{i}}=\frac{\sum_{i=1}^{n}Y_{i}Z_{i}}{n/2}-\frac{\sum_{i=1}^{n}Y_{i}(1-Z_{i})}{n/2}$, where the second equality holds under Assumption 1. Then we have,
$E_{\mathbf{Z}\sim\tilde{P}_{IGR}}[\hat{\tau}$] $=\frac{\sum_{i=1}^{n}E[Y_{i}Z_{i}]}{n/2}-\frac{E[Y_{i}(1-Z_{i})]}{n/2}$
$=\frac{\sum_{i=1}^{n}E[Y_{i}Z_{i}\mid Z_{i}=1]\tilde{P}_{IGR}(Z_{i}=1)+\sum_{i=1}^{n}E[Y_{i}Z_{i}\mid Z_{i}=0]\tilde{P}_{IGR}(Z_{i}=0)}{n/2}$-
$\quad\frac{\sum_{i=1}^{n}E[Y_{i}(Z_{i})\mid Z_{i}=1]\tilde{P}_{IGR}(Z_{i}=1)+\sum_{i=1}^{n}E[Y_{i}(1-Z_{i})\mid Z_{i}=0]\tilde{P}_{IGR}(Z_{i}=0)}{n/2}$
$=\frac{\sum_{i=1}^{n}Y_{i}(1)\cdot\frac{1}{2}}{n/2}-\frac{\sum_{i=1}^{n}Y_{i}(0)\cdot\frac{1}{2}}{n/2}\quad\quad\text{ by consistency and mirror property}$
$=\frac{1}{n}\sum_{i=1}^{n}Y_{i}(1)-Y_{i}(0)$
F.2 Bias in the Presence of Interference
When interference is present, a common effect estimand of interest is the global average treatment effect, $\tau_{GATE}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}(\mathbf{1})-Y_{i}(\mathbf{0})$. In general, designs that satisfy the mirror property and have equal-sized arms do not produce unbiased difference-in-means estimates of the global average treatment effect. Thus, while violation of the mirror property can introduce some bias, it may be “worth it”, so long as it sufficiently reduces interference bias.
F.2.1 Bias of an Unrestricted Design
Let $Y_{i}(Z_{i},\mathbf{Z}_{-i})$ denote the potential outcome under the assignment vector where individual $i$ is assigned to $Z_{i}$ and the remaining individuals to $\mathbf{Z}_{-i}.$ Consider an unrestricted design that satisfies the mirror property. Since $P(\mathbf{Z}=\mathbf{z})=P(\mathbf{Z}=1-\mathbf{z})$, it also holds that $P(Z_{i}=1)=P(1-Z_{i}=1)=\frac{1}{2}$. Then, under Assumption 1 we have,
$E[\hat{\tau}$] $=\sum_{i=1}^{n}\frac{E[Y_{i}Z_{i}]}{\frac{n}{2}}-\frac{E[Y_{i}(1-Z_{i})]}{\frac{n}{2}}$
$=\sum_{i=1}^{n}\frac{E[Y_{i}(1,\mathbf{Z}_{-i})]\cdot P(Z_{i}=1)}{\frac{n}{2}}-\frac{E[Y_{i}(0,\mathbf{Z}_{-i})]\cdot P(Z_{i}=0)}{\frac{n}{2}}$
$=\frac{1}{n}\sum_{i=1}^{n}E[Y_{i}(1,\mathbf{Z}_{-i})]-E[Y_{i}(0,\mathbf{Z}_{-i})]$
where the expected value is taken over the possible assignment vectors sampled from the design. The bias in the estimate is therefore
$\text{Bias}_{\text{unrestricted}}$ $=E[\hat{\tau}]-\tau_{GATE}$
$=\frac{1}{n}\sum_{i=1}^{n}E[Y_{i}(1,\mathbf{Z}_{-i})]-E[Y_{i}(0,\mathbf{Z}_{-i})]-\frac{1}{n}\sum_{i=1}^{n}Y_{i}(\mathbf{1})-Y_{i}(\mathbf{0})$ (F.1)
F.2.2 Bias of an IGR Design
Now, consider instead an IGR design with fitness function $f$ and restriction rule $r$, such that an assignment vector $\mathbf{z}$ is accepted if and only if $r(f(\mathbf{z}))=1$. Let $\mathcal{A}$ denote the event $r(f(\mathbf{z}))=1$. Suppose this design does not satisfy the mirror property, so $P(\mathbf{Z}=\mathbf{z})\neq P(\mathbf{Z}=1-\mathbf{z})$ and therefore $P(Z_{i}=1\mid\mathcal{A})$ may not be equal to $\frac{1}{2}$. Instead, for every $i$, there exists some perturbation $\Delta_{i}\in(-\frac{1}{2},\frac{1}{2})$ such that $P(Z_{i}=1\mid\mathcal{A})=\frac{1}{2}+\Delta_{i}$ and $P(Z_{i}=0\mid\mathcal{A})=\frac{1}{2}-\Delta_{i}$. We continue to operate under Assumption 1, so that the treatment and control arms are equally sized.
Proposition 2 (Zero-sum perturbation).
Under Assumption 1, $\sum_{i=1}^{n}\Delta_{i}=0$.
Proof.
Since $Z_{i}\in\{0,1\}$, we have $E[Z_{i}\mid\mathcal{A}]=P(Z_{i}=1\mid\mathcal{A})=\frac{1}{2}+\Delta_{i}$. Thus,
$\sum_{i=1}^{n}E[Z_{i}\mid\mathcal{A}]$ $=\sum_{i=1}^{n}\frac{1}{2}+\Delta_{i}$
$=\frac{n}{2}+\sum_{i=1}^{n}\Delta_{i}$
Pulling the expectation out of the sum, the left-hand side becomes $\sum_{i=1}^{n}E[Z_{i}\mid\mathcal{A}]=E[\sum_{i=1}^{n}Z_{i}\mid\mathcal{A}]=\frac{n}{2}$ under Assumption 1. Thus, $\sum_{i=1}^{n}\Delta_{i}=0$. ∎
Recall that in IGR, we enumerate a pool of candidate assignment vectors $\mathcal{Z}_{pool}$, then filter out those that do not meet design desiderata. We can write the difference-in-means estimate under an IGR design as $E_{\mathcal{Z}_{pool}}[\hat{\tau}\mid\mathcal{A}]$. Then, under Assumption 1 we have,
$E[\hat{\tau}\mid\mathcal{A}]$ $=\sum_{i=1}^{n}\frac{E[Y_{i}Z_{i}\mid\mathcal{A}]}{\frac{n}{2}}-\frac{E[Y_{i}(1-Z_{i})\mid\mathcal{A}]}{\frac{n}{2}}$
$=\sum_{i=1}^{n}\frac{E[Y_{i}(1,\mathbf{Z}_{-i})\mid\mathcal{A}]\cdot P(Z_{i}=1\mid\mathcal{A})}{\frac{n}{2}}-\frac{E[Y_{i}(0,\mathbf{Z}_{-i})\mid\mathcal{A}]\cdot P(Z_{i}=0\mid\mathcal{A})}{\frac{n}{2}}$
$=\sum_{i=1}^{n}\frac{E[Y_{i}(1,\mathbf{Z}_{-i})\mid\mathcal{A}]\cdot(\frac{1}{2}+\Delta_{i})}{\frac{n}{2}}-\frac{E[Y_{i}(0,\mathbf{Z}_{-i})\mid\mathcal{A}]\cdot(\frac{1}{2}-\Delta_{i})}{\frac{n}{2}}$
$=\frac{1}{n}\sum_{i=1}^{n}E[Y_{i}(1,\mathbf{Z}_{-i})\mid\mathcal{A}]-E[Y_{i}(0,\mathbf{Z}_{-i})\mid\mathcal{A}]+\frac{2}{n}\sum_{i=1}^{n}\Delta_{i}\left(E[Y_{i}(1,\mathbf{Z}_{-i})\mid\mathcal{A}]+E[Y_{i}(0,\mathbf{Z}_{-i})\mid\mathcal{A}]\right)$
The bias in the estimate is therefore
$\text{Bias}_{\text{IGR}}$ $=E[\hat{\tau}\mid\mathcal{A}]-\tau_{GATE}$
$=\frac{1}{n}\sum_{i=1}^{n}E[Y_{i}(1,\mathbf{Z}_{-i})\mid\mathcal{A}]-E[Y_{i}(0,\mathbf{Z}_{-i})\mid\mathcal{A}]$
$\quad\quad+\frac{2}{n}\sum_{i=1}^{n}\Delta_{i}\left(E[Y_{i}(1,\mathbf{Z}_{-i})\mid\mathcal{A}]+E[Y_{i}(0,\mathbf{Z}_{-i})\mid\mathcal{A}]\right)$
$\quad\quad-\frac{1}{n}\sum_{i=1}^{n}Y_{i}(\mathbf{1})-Y_{i}(\mathbf{0})$ (F.2)
IGR reduces absolute bias compared to an unrestricted design if $|\text{Bias}_{\text{IGR}}|<|\text{Bias}_{\text{unrestricted}}|$.
F.2.3 Bias of an Unrestricted vs IGR design for an Additive Outcome Model
When does the above inequality hold? To gain some insight, we look at a specific outcome model with an additive direct effect and an additive spillover effect on individuals assigned to the control arm.
Definition 1 (Additive outcome model).
$Y_{i}(\mathbf{Z})$ $=\alpha_{i}+\tau Z_{i}+\delta h(\mathbf{Z}_{-i})(1-Z_{i})$ (F.3)
where $\alpha_{i}$, $\tau$, $\delta$ are scalars and $h:\mathbb{R}^{N}\times\mathbb{R}^{N\times d}\mapsto\mathbb{R}_{\geq 0}$ is an exposure function that maps the treatment assignments of everyone other than $i$ to a non-negative scalar value. Assume $h(\mathbf{0})=0$.
With this outcome model, we have $\tau_{GATE}=\frac{1}{n}\sum_{i=1}^{n}\alpha_{i}+\tau-\alpha_{i}=\tau$. We also have $Y_{i}(1,\mathbf{Z}_{-i})=\alpha_{i}+\tau$ and $Y_{i}(0,\mathbf{Z}_{-i})=\alpha_{i}+\delta h(\mathbf{Z}_{-i})$. Thus, we can write $\text{Bias}_{\text{Unrestricted}}$ as
$\text{Bias}_{\text{Unrestricted}}$ $=\frac{1}{n}\sum_{i=1}^{n}E[\alpha_{i}+\tau]-E[\alpha_{i}+\delta h(\mathbf{Z_{-i}})]-\tau$
$=\frac{1}{n}\sum_{i=1}^{n}-\delta E[h(\mathbf{Z}_{-i})]$
[…]